OpenStudy (anonymous):

1=9x-(3x-8) I cannot seem to get this one. Please help!

OpenStudy (anonymous):

6x+8=1 6x=-7 x=-7/6

OpenStudy (anonymous):

expand the brackets as such: 1=9x-3x+8 now solve

OpenStudy (anonymous):

Easier than you think, the parentheses are a bit of a trick

OpenStudy (anonymous):

1=9x-3x+8 as heedcom mentioned then 1=6x+8 as 9x-3x=6x then bring 8 to the right hand side of the equals sign so now it's -8 and we have 1-8=6x

OpenStudy (anonymous):

1-8 = -7 so we have: -7=6x OR 6x= -7 where Aztek left of. GL

OpenStudy (anonymous):

I get it now. Thank you so much!